In market A, a firm with market power faces an inverse demand curve of P = 10 – Q and a marginal cost that is constant at $2. In market B, a firm with market power faces an inverse demand curve of P = 8 – 0.75Q and a marginal cost of $2. Producer surplus in market A is _____ than in market B.

$4 higher=correct

how?
Answer #1

Answer
the firm produces at MR=MC
where
MR=10-2Q ....... the MR curve is double sloped than the demand curve
equating MR=MC
10-2Q=2
Q=8/2=4
P=10-4=6
PS=(P-MC)*Q=(6-2)*4=$16
------
MR=8-1.5Q
equating MR=MC

8-1.5Q=2
1.5Q=6
Q=4
P=8-0.75*4=5
PS=(5-2)*4=$12
-----
the difference =16-12=$4

$4 higher